OpenStudy (anonymous):

How long will it take for the value of an account to be $890 if $350 is deposited at 11% interest compounded continously? Thanks!! ;)

OpenStudy (anonymous):

Use the rule of 72. 72/ 11 = 6.5. 890 is 2.5 times 350 so with the answers you have it would be 8.48 yrs.

OpenStudy (anonymous):

8.48 is the answer

OpenStudy (mathstudent55):

A = pe^(rt) where p = principal amount (what you start with) e = base of natural log r = interest rate (as a decimal: 5% = 0.05) t = time (number of years) A = amount at end of t years at r interest

OpenStudy (anonymous):

@blast234 , yes it is the answer.

OpenStudy (anonymous):

Welcome

OpenStudy (anonymous):

But I would like to understand it better, in this case what will be the time...???

OpenStudy (anonymous):

Oh, then its Ok.

OpenStudy (anonymous):

Mathstudent was also helpful XD

OpenStudy (anonymous):

he deserves a medal too

OpenStudy (mathstudent55):

A = pe^(rt) 890 = 350e^(0.11t) e^(0.11t) = 890/350 = 89/35 ln(e^(0.11t)) = ln (89/35) 0.11t = ln (89/35) t = (ln (89/35))/0.11

OpenStudy (anonymous):

Thank you so much!!! You deserved it =D

OpenStudy (mathstudent55):

wlcm